1. If the number of employed is 290,000, the civilian labor force is 320,000, the number of discouraged workers is 35,000, the unemployment rate is:

A. 14.7%

B. 9.4%

C. 12%

D. 8.8%

E. 19.1%

2. Assume an economy with the following info:

C=$275,000 + 0.6Yd

I= 215,000

G= 190,000

X= 209,000

IM= 0.1Y

T=100,000

Question:

a. Calculate national saving at Y*

b. With potential GDP at $3,000,000. How much should the government increase/decrease its spending? How much should the government increase/decrease its taxes?
Answer #1

Answer-1 The correct option is B.) 9.4%

The unemployment rate is the number of unemployed people divided by the total number of people in the civilian labor force.

Civilian labor force includes the unemployed plus employed people and no one else.

Number of unemployed people = 320,000-290,000

= 30,000

Unemployment rate = 30,000/320,000

= 0.09375

Hence as a percentage it is 9.4%

Answer-2

Y = Ct It G + NX YE 275,000+ 0.6CY - 100,000) + 215,000 + 190,000+ Y-0.6Y+0.14 = 275,000 - 60,000 + 215,000 + 190,000 + 209,0Government should increase its spending by $536,800 - Tax multiplier - MPC - IMPL - 0.6 (2.50 : -1.5 - AT= (1,342,000) - $894

Y = Ct It G + NX YE 275,000+ 0.6CY - 100,000) + 215,000 + 190,000+ Y-0.6Y+0.14 = 275,000 - 60,000 + 215,000 + 190,000 + 209,00 0.54= 829,000 V = 1658,000 no+(209,000 - 014) 1209,000 A.) National Saving S I = I + NX. = 215,000 + (209,000 - 163800) 2 $ 258,200 b.) Potential app = $3,000,000 Actual GDP = 1,658,000 MPC(= 0.6 Government Expenditure Multiplier : 1/1-MPC tooth ds 0.4 AY = 2.5.DG DG = (3,000,000) – (1,658,000) 2.5 a 1,342,000 2.5 2536800

Government should increase its spending by $536,800 - Tax multiplier - MPC - IMPL - 0.6 (2.50 : -1.5 - AT= (1,342,000) - $894.666.67 Government should decrease taxes by $894, 666.67